# A note on the electrostatic Born–Infeld equation with radial charge density

### 06 Apr 2026 v3

@@ 3.3 Vanishing ADM mass and spacelike hypersurfaces in $\mathbb{L}^{n+1}$
    if $q>{n\over 2}$, then $m_{\text{ADM}}(g)=0$, and hence, $(\mathbb{R}^{n},g,k)$ is isometric to an AF spakelike hypersurface of $\mathbb{L}^{n+1}$ as long as $(q,n)$ fulfills (2.8). In particular, Theorem 1.1 holds.
@@
We recall that the radial functions $\rho$ and $\varphi$ in the definition (3.13) of $(g,k)$ solves the conformal equations (3.1). By Proposition 3.1, it follows that $(\rho,\varphi)$ must satisfy Identity (3.2). Then, provided that $|\rho|\sim cr^{-q}$, this identity gives us
@@
By L’Hôpital’s rule, it follows that
$\lim_{r\to+\infty}\Big({\varphi^{\prime}\over r^{-2q+1}}\Big)=\Bigg(\lim_{r\to+\infty}{\sqrt{r^{2n-1}\varphi^{\prime}}\over r^{n-q}}\Bigg)^{2}=\Bigg(\lim_{r\to+\infty}\Bigg|{\big(\sqrt{r^{2n-1}\varphi^{\prime}}\big)^{\prime}\over(r^{n-q})^{\prime}}\Bigg|\Bigg)^{2}={c\sqrt{n-2}\over 2(n-q)}.$ (3.18)
@@
$m_{\text{ADM}}(g)=-{c\sqrt{n-2}\over 4(n-q)}\lim\limits_{r\to+\infty}r^{n-2q}.$
@@
From this fact, and using the Spacetime PET (which is needed only for case (iii)), we conclude cases (i–iii). The proof is completed. ∎
4 Further discussion
@@
We have established the existence of classical solutions to the equation (1.1) when $\rho$ is radial by using the conformal method and the Spacetime PET. For the future study of this approach, we make the following remarks:
@@
- Sharpness of the Spacetime PET. Regarding the ADM mass, we emphasize that the appearance of negative mass in Proposition 3.3 does not contradict the Spacetime PET. This is because the decay rate of $k$ at infinity in the first two cases of the proposition is either critical or subcritical relative to the decay assumptions imposed on the second fundamental form in the theorem. Combined with the construction of Chruściel [4], which provides an example of vacuum constraint solutions with negative mass in the case where $g-\delta_{\text{Euc}}$ decays subcritically at infinity, these results demonstrate that the decay conditions on both $g$ and $k$ stated in the theorem are sharp.
@@
- The Birkhoff Theorem Point of View. Proposition 3.3 can be strengthened to state that, in all three cases, $(\mathbb{R}^{n},g,k)$ is isometric to a spacelike hypersurface of Lorentz–Minkowski spacetime, regardless of the mass. Consequently, Theorem 1.1 remains valid under the weaker assumption $q>2$ in all dimensions. In fact, since $\rho$ and $\varphi$ are radial, $(g,k)$ is, by definition, regular and spherically symmetric. Then, by Birkhoff’s theorem, $(\mathbb{R}^{n},g,k)$ is automatically an asymptotically flat (AF) initial data set in $\mathbb{L}^{n+1}$, without any need to consider its mass. Nevertheless, as noted in the Introduction, we prefer the current statement, since it seems to apply to a broader class of $\rho$ beyond the radial case.
@@
- Lower regularity on charge density. In physics, it is important to study $\rho$ under Sobolev regularity, for instance, $\rho$ may be a finite superposition of point charges, that is, $\rho=\Sigma_{i=1}^{k}a_{i}\delta_{x_{i}}$, where $a_{i}\in\mathbb{R}$ and $\delta_{x_{i}}$ is the Dirac delta distribution centered at $x_{i}\in\mathbb{R}^{n}$. In our setting, Sobolev regularity does not pose any difficulty at the stage of applying the conformal method to construct initial data sets with prescribed mean curvature. However, Hölder regularity is required in order to satisfy the assumptions in the results of Chruściel–Maerten [5] and Eichmair [7], which are used to establish the spacetime PET. Therefore, if the regularity assumptions in the spacetime PET could be weakened, the assumptions in Theorem 1.1 would immediately become less restrictive as well. This appears to be a feasible direction, inspired by the work of Lee and LeFloch [8], who showed that the Positive Mass Theorem continues to hold for metrics with only Sobolev regularity, where curvature is defined merely in a distributional sense. If a similar result can be extended to the spacetime PET, then Theorem 1.1 could be further strengthened by assuming only Sobolev regularity on $\rho$.
Taken together, these remarks indicate possible directions for extending our results and suggest that the framework developed here might be adaptable to more general settings.

### 16 Apr 2026 v5

@@ 3.3 Vanishing ADM mass and spacelike hypersurfaces in $\mathbb{L}^{n+1}$
    if $q>{n\over 2}$, then $m_{\text{ADM}}(g)=0$, and hence, $(\mathbb{R}^{n},g,k)$ is isometric to an AF spacelike hypersurface of $\mathbb{L}^{n+1}$ as long as $(q,n)$ fulfills (2.8). In particular, Theorem 1.1 holds.
@@
We recall that the radial functions $\rho$ and $\varphi$ in the definition (3.13) of $(g,k)$ solves the conformal equations (3.1). On one hand, by Proposition 3.1, it follows that $\varphi^{\prime}\geq 0$ and $(\rho,\varphi)$ must satisfy Identity (3.2). Then, provided that $|\rho|\sim cr^{-q}$, this identity gives us
@@
Applying L’Hôpital’s rule to the ratio above, it follows that
$\lim_{r\to+\infty}\Big({\varphi^{\prime}\over r^{-2q+1}}\Big)=\Bigg(\lim_{r\to+\infty}{\sqrt{r^{2n-1}\varphi^{\prime}}\over r^{n-q}}\Bigg)^{2}=\Bigg(\lim_{r\to+\infty}\Bigg|{\big(\sqrt{r^{2n-1}\varphi^{\prime}}\big)^{\prime}\over(r^{n-q})^{\prime}}\Bigg|\Bigg)^{2}={c^{2}(n-2)\over 4(n-q)^{2}}.$ (3.18)
@@
$m_{\text{ADM}}(g)=-{c^{2}(n-1)\over 8(n-q)^{2}}\lim\limits_{r\to+\infty}r^{n-2q}.$
@@
From this fact, together with the Spacetime PET (which is needed only for case (iii)), we conclude cases (i–iii). The proof is completed. ∎
@@
We have established the existence of classical solutions to the equation (1.1) when $\rho$ is radial by using the conformal method and the Spacetime PET. To conclude the article and provide further discussion of the approach, we make the following remarks.
@@
Regarding the sign of mass, we emphasize that the appearance of negative mass in Proposition 3.3 does not contradict the Spacetime PET. This is because the decay rate of $k$ at infinity in the first two cases of the proposition is either critical or subcritical relative to the decay assumptions imposed on the second fundamental form in the theorem. Combined with the construction of Chruściel [4], which provides an example of vacuum constraint solutions with negative mass in the case where $g-\delta_{\text{Euc}}$ decays subcritically at infinity, these results demonstrate that the decay conditions on both $g$ and $k$ stated in the theorem are sharp.
@@
It is also worth noting that Proposition 3.3 can be strengthened to state that, in all three cases, $(\mathbb{R}^{n},g,k)$ is isometric to a spacelike hypersurface of Lorentz–Minkowski spacetime, independently of the value of the mass. Consequently, Theorem 1.1 remains valid under the weaker assumption $q>2$ in all dimensions. In fact, since $\rho$ and $\varphi$ are radial, $(g,k)$ is, by definition, regular and spherically symmetric. Then, it follows by Birkhoff’s theorem that $(\mathbb{R}^{n},g,k)$ is automatically an AF initial data set in $\mathbb{L}^{n+1}$, without any need to consider its mass. Nevertheless, as noted in the Introduction, we prefer the current statement, since it seems to apply to a broader class of $\rho$ beyond the radial case.
@@
Finally, we remark that solutions to the electrostatic Born–Infeld (1.1) in our result are classical and spacelike, representing an improvement in solvability compared to those in [2, 3]. Conversely, the regularity assumption on $\rho$ in our result is a bit stronger than in existing works, where $\rho$ is only required to be Sobolev regular. More precisely, Sobolev regularity does not present any difficulty at the stage of applying the conformal method to construct initial data sets with prescribed mean curvature, but Hölder regularity for $\rho$ is needed in order to satisfy the assumptions in the spacetime PET version of Chruściel–Maerten [5] and Eichmair [7]. Therefore, if the regularity assumptions in the spacetime PET could be weakened, the assumptions in Theorem 1.1 would immediately become less restrictive as well. This appears to be a feasible direction, inspired by the work of Lee and LeFloch [8], who showed that the Positive Mass Theorem continues to hold for metrics with only Sobolev regularity. If a similar result can be extended to the spacetime PET, then Theorem 1.1 could be further strengthened by assuming only Sobolev regularity on $\rho$.